OpenStudy (anonymous):

What is 5.8 × 106 mg/s in grams per minute? A. 9.7 × 109 g/min B. 9.7 × 107 g/min C. 3.48 × 105 g/min D. 3.48 × 103 g/min

OpenStudy (sjg13e):

C. \[? g/m =\frac{ 5.8x10^{6}mg}{ s }|\frac{ 1g }{ 1000mg }|\frac{ 60s }{ 1m }\] I do the train and caboose method...

OpenStudy (sjg13e):

Just cross multiply and then divide. Answer: 3.48 x 10^5

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

What is 5.6 × 10−3 yd/year in feet per day? A. 6.813 × 10−7 ft/day B. 6.132 × 10−6 ft/day C. 5 × 10−6 ft/day D. 4.6 × 10−5 ft/day

OpenStudy (anonymous):

can someone please